OpenStudy (anonymous):

Solving Equations with multiplying and dividing. 1. -y=-32 2. 7r=-56 3. t/-3 =12 4. 4= s/-14 5.b/47=-2 P.S. "The / means its a fraction."

OpenStudy (anonymous):

anyone can help me?

OpenStudy (keigh2015):

I don't know this I am sorry. @alexistheking777 come take a look at this and @ianmack904 can you come look at this?

OpenStudy (anonymous):

Its fine i been looking for help all day.

OpenStudy (keigh2015):

I am sorry I couldn't be of anymore help

OpenStudy (anonymous):

dont be srry, u ii

sammixboo (sammixboo):

Let's do #5 *-* We have \(\tt \color{red}{\dfrac{b}{47}= -2}\) To solve this you must: \(\tt \color{red}{Multiply~47~on~both~sides~of~the~equation.}\) So, now we have \(\tt \color{red}{\dfrac{b}{47}\times47= -2\times47}\) Do you know how to simplify this? Once you do, you will get your answer :)

OpenStudy (anonymous):

so do -2x47?

sammixboo (sammixboo):

Right

sammixboo (sammixboo):

And that's your answer for #5. Tell me what you got :)

OpenStudy (anonymous):

b=(-94)

sammixboo (sammixboo):

Correcto *-* Now let's do #4: \(\tt \color{red}{4= \dfrac{s}{-14}}\) Do you have an idea on how to solve this?

OpenStudy (anonymous):

i would do -14/4?

sammixboo (sammixboo):

Not quite, you would \(\tt \color{red}{multiply}\), since you are originally \(\tt \color{red}{dividing}\).

sammixboo (sammixboo):

So, \(\tt \color{red}{-14\times4}\)

OpenStudy (anonymous):

-56

OpenStudy (anonymous):

s=(-56)

sammixboo (sammixboo):

Correct again ^-^ #4 is -56. Let's do #3: \(\tt \color{red}{\dfrac{t}{-3}=12}\) Do you know what to do on this one?

OpenStudy (anonymous):

12x-3?

sammixboo (sammixboo):

Right, which is?

OpenStudy (anonymous):

t=-36

sammixboo (sammixboo):

Right *-*, so #3 is -36 Now for #2, we do the opposite since we are originally multiplying. \(\tt \color{red}{7r=-56}\) We have to \(\tt \color{red}{divide~7~on~both~sides}\). What do you think the answer is?

OpenStudy (anonymous):

SO....-56/7

sammixboo (sammixboo):

Yeppers :)

OpenStudy (anonymous):

UM...I GOT SUM REALLY LONG NUMBER

sammixboo (sammixboo):

You shouldn't. Make sure you type -56/7 and not 7/-56

OpenStudy (anonymous):

i got -8.142857143

sammixboo (sammixboo):

Sure you typed everything correct?

OpenStudy (anonymous):

yep stll got the long number

sammixboo (sammixboo):

You typed -57/7, not -56/7

OpenStudy (anonymous):

X___x

OpenStudy (anonymous):

I GOT -8

sammixboo (sammixboo):

^-^ It's ok. It happens a lot, and that is right, so #2 is -8

OpenStudy (anonymous):

r=-8

OpenStudy (anonymous):

i think ik #1

sammixboo (sammixboo):

What do you think it is?

OpenStudy (anonymous):

y=-32

OpenStudy (anonymous):

cause u remeber the teachers saying if the letter by itself it counts as a 1 in front of it

sammixboo (sammixboo):

Close, but not quite. Now for #1, we have \(\tt \color{red}{-y = -32}\), since they are both negative, your answer should be positive

OpenStudy (anonymous):

OH OK

OpenStudy (anonymous):

y=32

sammixboo (sammixboo):

If y = -32, then we would have - - 32 = -32, which isn't right :P

sammixboo (sammixboo):

Right *-*

sammixboo (sammixboo):

Good job :)

OpenStudy (anonymous):

YAY...i feel smart

OpenStudy (anonymous):

i think i can do the other 16 by myself

sammixboo (sammixboo):

Okie dokie. Once you finish you can post it here and I can check it if you want.

OpenStudy (anonymous):

ok